please show work
-78.33 POINTS SCALCET8 3.4.062. If h(x) = 9(4) = 5 + 4f(x), where f(4) = 5 and f(4) = 2, find h(4). Show My Work (Required
Answer #1

h(x) -JS+५६(५) (4): २ , fu) 5. ____b4) - १ (0+ 4 1 x २J544fox) (x) nix) = f (0) stutxi (4) = a + (4) J5+ Lf(५) 2 x २ J5 + 405